An airplane flying headwind travels the 1800 mile flying distance between two cities in 3 hours and 36 minutes. One the return flight, the distance is traveled in 3 hours. Find the air speed of the plane and the speed of the wind, assuming that both remain constant.
Answer by stanbon(75887)

An airplane flying headwind travels the 1800 mile flying distance between two cities in 3 hours and 36 minutes. One the return flight, the distance is traveled in 3 hours. Find the air speed of the plane and the speed of the wind, assuming that both remain constant.
-------------------------
Against wind DATA:
distance = 1800 miles; time = (216/60) hrs ; rate = d/r = 500 mph
-----
With wind DATA:
distance = 1800 miles ; time = 3 hrs ; rate = d/r = 600 mph
---------------------
Equations:
p + w = 600
p - w = 500
---
2p = 1100
p = 550 (speed of the plane in still air
---
p + w = 600
w = 50 mph (speed of the wind current)
==========================================
Cheers,
Stan H.